determine whether or not the following given ordered pairs, labeled a,b and c are solutions to the given equations.
y=3x-1 A.(1,2)
B.(-1,-4)
C.(2,3)
how do I do this?

Answer by stanbon(75887)

determine whether or not the following given ordered pairs, labeled a,b and c are solutions to the given equations.
y=3x-1
A.(1,2)
2 = 3(1)-1
2 = 3-1
2 = 2

Yes
-------------------
B.(-1,-4)
-4 = 3(-1)-1
-4 = -3-1
-4=-4
Yes
-----------------
C.(2,3)
3 = 3(2)-1
3 = 6-1
3 = 5
No
-----------------
Cheers,
Stan H.